Distinguish between GDP deflator and Consumer Price Index (CPI). (7)

(b) The CPI of Namibia was 101 in December 2018. It then rose to 133 in December 2019,

(i) Determine the annual rate of inflation over the period. (4)

(ii) Explain any shortcoming of real GDP from a welfare perspective (7)

(iii) Distinguish between nominal GDP and real GDP, and explain why nominal GDP growth of

an economy is usually higher than real growth.


1
Expert's answer
2021-04-29T10:21:18-0400

Solution:

a.). The differences between GDP deflator and Consumer Price Index (CPI) include the following:

·        The GDP deflator measures the prices of all products and services produced, whereas the CPI measures the prices of only the products and services purchased by consumers.

·        The GDP deflator only includes the products produced domestically, while CPI includes both products imported and produced domestically.

·        The GDP deflator assigns changing weights on the prices of multiple products, whereas CPI assigns fixed weights.

 

b.). i). The annual rate of inflation =


"\frac{Current \;CPI - Past\; CPI}{Current \;CPI} \times 100 = \frac{133 - 101}{133} \times 100 = 24.1\%"


The annual rate of inflation = 24.1%

 

ii). The limitations of real GDP from a welfare perspective include the following:

·        It only counts goods that pass through official channels and excludes non-market transactions such as home production and black-market activity.

·        It does not make adjustments for leisure time.

·        It does not account for or represent the degree of income inequality in society.

 

iii). Nominal GDP is a method of measuring the value of all products and services produced by an economy at current market prices in a financial year. No inflation is adjusted in Nominal GDP.

On the other hand, Real GDP is a method of measuring a country’s output in terms of the value of its products and services using the base year prices in a financial year. Real GDP takes into account adjustments for inflation changes.

 

The nominal GDP growth of an economy is usually higher than real growth since it is not adjusted for inflation. Real GDP is adjusted for inflation which reduces its value.